# TE-NeXt: A LiDAR-Based 3D Sparse Convolutional Network for Traversability Estimation

### v2

@@ Abstract
Among the multitude of scenarios in which autonomous robots are intended to operate, natural environments present the most significant challenges in the context of traversability estimation. Conversely, although comparatively less complex, structured environments are the most prevalent in the broader sense of autonomous navigation due to the abundance of human-designed spaces, which offer more predictable features for navigation systems. Therefore, both of these considerations lead to the conclusion that there is a need to develop methods that are specifically tailored to unstructured environments, while ensuring that the performance in structured environments is not significantly compromised. Existing methods are vulnerable to the influence of stationary environmental conditions or are constrained by the platform or vehicle on which they are developed. In this context, this paper makes a contribution to the state of the art with TE-NeXt. This network is a customised and efficient architecture for traversability estimation from sparse LiDAR point clouds based on a encoder-decoder topology that includes several modifications regarding: (i) the input features; (ii) the structure of encoder-decoder resolution levels; and (iii) the constitution of the 3D convolutional block. The modifications provide the robot with a deeper understanding of traversability in natural and unstructured environments. Thus, the designed architecture presents a trade-off solution, exhibiting superior performance in unstructured environments (the Rellis-3D dataset) while maintaining high reliability and robustness in urban environments (the SemanticKITTI and the SemanticUSL dataset). The implementation of the presented approach is available at an open repository to ensure result reproducibility.
Index Terms:
Autonomous mobile robots, LiDAR, Traversability Estimation, 3D Sparse Convolution, Semantic Segmentation.
I Introduction
@@
Figure 1: Point cloud-based traversability estimation. TeNeXt performs inferences about the different instances that conform the map. Each point cloud represents a pose within the 3D map of the environment. Areas identified as potentially traversable are shown in   and non-traversable areas are shown in  .
@@
To enhance the efficacy of specialised techniques for off-road traversability estimation, this study proposes TE-NeXt, a LiDAR-based 3D sparse convolutional architecture capable of discerning traversable and non-traversable areas captured by a LiDAR sensor (Figure 1), enriching the perception of the environment with semantic information in addition to geometric information. In summary, the main contributions of this paper are:
@@
  • Applicability of the proposed method in an autonomous navigation system in dynamic environments.
@@
The remaining sections of the paper are structured as follows: Section II outlines the prevailing methodologies employed in addressing the issue of traversability estimation. Section III presents the proposed architecture in detail. Following, Section IV describes an exhaustive set of experiments carried out on different datasets with the aim of comparing the current approach with other cutting-edge methods.
II RELATED WORK
@@
II-A Vision-based traversability estimation
@@
II-B Geometric and hybrid methods for traversability estimation
@@
Terrain classification into road and off-road
categories.
@@
Schreiber et al., 2024 [35] RGB images Mean Teacher-CNN Method of relative traversability estimation that uses sparsely annotated point pairs. No
@@
II-C Self-supervised traversability estimation
@@
II-D Current proposal
This paper presents TE-NeXt, an encoder-decoder architecture optimised for traversability estimation in unstructured environments. The approach presented here can be classified as a supervised method, in accordance with the paradigm described in Subsection II-B, as it is a supervised paradigm method. However, it employs local point clouds, a distinctive feature that sets it apart from other methods, such as those described in [23, 24], which utilise point cloud maps and preprocess the data. Moreover, in contrast to the techniques outlined in Subsection II-A, TE-NeXt employs an optical sensor that is invariant to changes in light conditions (LiDAR sensor). Furthermore, the proposed methodology is designed to overcome the limitations associated with self-supervised learning approaches. Unlike existing methods, it does not require a specific platform, making it a more general and secure solution for all types of vehicles. Table I outlines the distinctive features of the proposed approach in comparison to existing, state-of-the-art methodologies.
@@ II-D Current proposal
III TE-NeXt: AN ARCHITECTURE FOR TRAVERSABILITY ESTIMATION
@@
III-A Network Architecture Overview
@@
III-B TE-NeXt residual block
@@ IV-A Datasets
  • SemanticUSL [47]: Comprises 16,578 unlabeled LiDAR scans, which are intended for unsupervised domain adaptation training, and an additional 1,200 labeled scans for evaluation purposes. SemanticUSL adopts the same data format and ontology as the well-established SemanticKITTI dataset, making it convenient for researchers to utilize in their domain adaptation studies. The data collection process for SemanticUSL involved the use of a Clearpath Warthog robotics platform equipped with an Ouster OS1-64 LiDAR sensor.
@@
IV-B Training and evaluation process
Given the considerable discrepancy in the quantity of labelled data available from urban and natural environments, being the latter the primary focus of this research, the objective was to achieve a more balanced distribution of data across these domains. In particular, 10,700 clouds of unstructured environments have been selected, belonging to four of the five sequences in Rellis-3D. Additionally, the initial four SemanticKITTI sequences have been selected, thus ensuring 11,100 clouds of structured environments. Furthermore, the validation and testing process have been conducted with sequences from both datasets, incorporating all the SemanticUSL data to assess the degree of generalisation of the neural network.
Table II specifies the hyperparameters and settings used to train the model from randomly initialized weights. The optimizer used AdamW with an initial Learning Rate (LR) of 5e-3 along to Cosine Annealing scheduler causes the LR to follow a cosine function over the same training epoch. It performs periodic “warm restarts” where the learning rate is reset back to the initial value after a certain number of batches. This simulates the restarting of the learning process using the weights from the previous run as a good starting point. The periodic restarts help the model escape from local minima and saddle points. Since the model has been trained for a binary segmentation task, the F1 score is a suitable metric for imbalanced datasets and considers both: precision (the proportion of true positive instances among all positive predictions) and recall (the proportion of true positive predictions among all actual positive instances). This guarantees that the model is not optimising for one metric at the expense of the other. Moreover, should the F1 score on the validation set cease to improve or begin to decline, training can be terminated prematurely to avert overfitting.
TABLE II: Hyperparameters selected in training process
Scratch
@@ IV-B Training and evaluation process
Figure 4: Design process of the proposed architecture, TE-NeXt, starting from a base MinkU-Net34C. All experiments are detailed in Table III.
IV-C Ablation study
@@
IV-C1 Input data
@@
  • Features: In contrast to 2D convolutions, where the input and output tensors have a fixed number of feature dimensions, typically called channels (e.g. RGB channels for images), 3D sparse convolutions allow sparse tensors to have an arbitrary number of feature dimensions. Each non-zero element in the sparse tensor can have a feature vector of arbitrary length. Therefore, a study of the most discriminating features for estimating traversability was carried out. Among the most representative are those shown in Figure 4 (experiments I.4, I.5, and I.6, referred to in Table III). Nevertheless, the vector of 1’s (initialising the feature associated with each point to a value of 1) was selected as the primary candidate for future modifications, as it is equally discriminant and, at the time of inference, represents an improvement in efficiency.
IV-C2 Global Design
  • Cardinality: This term was introduced by ResNext [48], and refers to the number of parallel branches within each block. Related to this, a number of different configurations have been assessed. The initial configuration employed was the MinkU-Net34 network, comprising (2, 3, 4, 6, 2, 2, 2, 2, 2) branches per block. In order to reduce the number of model parameters, it was decided to test the MinkU-Net18 and MinkU-Net14, experiments G.1 and G.2. The smaller versions, in particular MinkU-Net14, exhibit a notable reduction in the number of branches per convolutional block in comparison to MinkU-Net34. This leads to a decrease in the number of filters and, consequently, in the number of trainable parameters in the model. In conclusion, through a meticulous examination of smaller architectural configurations, highly efficient MinkU-Net configurations have been identified that strike a balance between model size and performance. This illustrates that it is frequently feasible to significantly reduce the complexity of neural networks without compromising their capacity to learn and generalise. While this modification does not lead to an significant enhancement in performance, it does contribute to greater efficiency, as will be demonstrated in Table VII.
  • Skip connections: Unlike the U-Net architecture [37], TeNext presents a variation in the connections that are made between the respective parts of the encoder and decoder. In addition to the four skip connections that were previously implemented in the experiments, two and three connections corresponding to experiments G.3 and G.4 have been evaluated. Notwithstanding the fact that both exhibited a reduction of 2% in terms of the F1 metric, the more robust performance options, experiment G.2 (later mentioned as the alternative experimental branch) and experiment G.4 (a configuration comprising just two connections), were subjected to further investigation in subsequent experiments. This has been influenced by concepts from related research [23, 24].
@@ IV-C2 Global Design
TABLE III: Experimental modifications carried out during the design of the proposed architecture, for which performance is presented in Figure 4
@@
IV-C3 Micro Design
@@
  • Kernel size: Conventional CNNs typically utilise relatively small kernels. In contrast, Transformer-based architectures and CNN designs inspired by the Transformer often employ significantly larger kernels, with sizes beginning at 7x7 and beyond. This allows them to capture longer-range dependencies and broader context, contributing to their strong performance on various vision tasks. To explore the solutions proposed in Transformer-based architectures, a study of the convolution size has been conducted. Experiments M.3, M.4 and M.5, depicted in Figure 4 and detailed in Table III, demonstrate that a larger convolution window size leads to superior outcomes compared to the conventional convolution window sizes typically employed in convolutional networks (2, 3 or 5). Consequently, this variation results in an increase of 0.9% in the F1 metric with respect to the previous optimal result.
@@ IV-D Results
This section offers a comprehensive performance evaluation of the proposed method, comparing it with established semantic segmentation and traversability estimation models from the literature. To ensure a fair comparison, we focus on accessible semantic segmentation models and methods originally designed for traversability estimation. The analysis covers a diverse range of approaches, including multi-layer perceptrons, traditional 2D convolutional networks, and cutting-edge sparse convolutions for efficient 3D data processing. As emphasised throughout this paper, TE-NeXt is an architecture optimised for unstructured environments. However, it can also be regarded as a trade-off solution in urban environments. In order to facilitate a comprehensive and structured comparison, this analysis will be divided into three distinct sections, each of which will focus on a specific type of environment.
IV-D1 Performance in structured environments
@@
Table IV presents the results of several semantic segmentation methods for the estimation of traversability in urban environments, using the validation sequence of the SemanticKITTI dataset.
@@
In the context of traversability estimation, our proposed solution, while effective, is outperformed by methods like SqueezeSegv3 and SalsaNext when assessed in a 2D domain using image projections of 3D data (rows 3 and 5 of Table IV respectively). However, this 2D evaluation process introduces significant challenges and limitations. The projection from three-dimensional (3D) to two-dimensional (2D) space inevitably leads to information loss due to occlusions and depth ambiguities, where multiple points may map to a single pixel. When projecting segmentation results back to 3D space, issues such as one-to-many mapping and interpolation errors at object boundaries emerge, which result in a noticeable decrease in performance metrics [51, 52].
@@
Thus, these methods have also been evaluated in a 3D context in order to compare their performance under the same conditions as the other methods (rows 4 and 6 of Table IV respectively). A decrease in their performance is observed, which confirms the limitations of 2D to 3D projections. To illustrate, SqueezeSegv3 demonstrates a decrease in its F1 score from 96.5% in 2D to 95.8% in 3D, whereas SalsaNext exhibits a reduction from 95.9% to 94.9%. Consequently, while these methods may show superior results in 2D evaluations, they may not be optimal for critical traversability tasks in 3D real-world environments. In contrast, TE-NeXt demonstrates a solid balance between accuracy, precision, and reliability, achieving an F1 score of 96.0% in structured environments. Thus, making it more suitable and robust for semantic segmentation in challenging, 3D real-world scenarios.
@@
The comparison presented in Table IV has been performed by training the semantic segmentation models from scratch, with the exception of Cyl3D, PVKD and P-SVM, which were adapted to the traversability problem in [25] using the same approach. Furthermore, it has included additional traversability estimation methods, such as ProtoNet [53], MPTI [54] and PBG [20]. These metrics have been collected from the manuscripts of the aforementioned approaches, which have been subjected to the same testing conditions and code is not accessible.
TABLE IV: Traversability estimation performance of semantic segmentation methods in urban environments.
@@
TABLE V: Traversability estimation performance of semantic segmentation methods in mixed environments.
Mixed envs. Sem seg SemanticUSL
Methods 2D 3D F1 score Accuracy mIOU TPR TNR
PointNet [55] $\checkmark$ 90.2 90.1 82.0 94.9 85.7
PointNet++ [56] $\checkmark$ 87.1 87.0 77.1 91.4 83.1
SqueezeSegv3 [57] $\checkmark$ 94.6 97.1 93.0 94.2 98.2
SqueezeSegv3 [57] $\checkmark$ 93.8 94.3 89.1 91.6 96.7
SalsaNext [52] $\checkmark$ 95.0 97.3 93.5 95.9 97.8
SalsaNext [52] $\checkmark$ 94.0 94.7 90.2 94.4 97.8
MinkU-Net[39] $\checkmark$ 94.3 94.4 89.3 98.1 91.0
MinkU-Net[39]+ ConvNeXt [10] $\checkmark$ 95.0 95.1 90.6 98.3 92.2
TE-NeXt (ours) $\checkmark$ 95.0 95.0 90.6 97.9 92.5
IV-D2 Performance in mixed environments
Table V presents the results of various semantic segmentation methods for the estimation of traversability in outdoor environments. Contrarily to previous tests, this is a fully employed dataset for testing purposes, considering semi-structured environments.
Regarding the performance of the proposed model, it is shown that it is comparable with the best methods in terms of F1 score, even though all the modifications made with a bias towards performance in natural environments. Methods such as SqueezeSegV3 and SalsaNext show very competitive performance, especially in terms of accuracy and TNR. As with structured environments, these methods demonstrate inferior performance in three-dimensional contexts.
The proposed TE-NeXt model demonstrates impressive performance, achieving an F1 score and accuracy of 95.0%, along with a high mIoU of 90.6%. Moreover, TE-NeXt exhibits a remarkable ability to accurately identify traversable regions, as evidenced by its high true positive rate of 97.9%. Simultaneously, it proves high effectivity in excluding non-traversable areas, boasting a true negative rate of 92.5%. These results showcase the model’s robustness and reliability in distinguishing between traversable and non-traversable terrain.
IV-D3 Performance in natural environments
Table VI presents the results of various semantic segmentation methods for traversability estimation in natural environments using the Rellis-3D dataset. The comparison presented in Table VI is of particular relevance, given that natural environments represent the primary focus of the research. Unlike in urban environments, it is remarkable the good performance of the most classic method, PointNet, outperforming more modern methods in most of the metrics extracted.

### v3

@@ Abstract
Among the multitude of scenarios in which autonomous robots are intended to operate, natural environments present the most significant challenges in the context of traversability estimation. Conversely, although comparatively less complex, structured environments are the most prevalent in the broader sense of autonomous navigation due to the abundance of human-designed spaces, which offer more predictable features for navigation systems. Therefore, both of these considerations lead to the conclusion that there is a need to develop methods that are specifically tailored to unstructured environments, while ensuring that the performance in structured environments is not significantly compromised. Existing methods are vulnerable to the influence of stationary environmental conditions or are constrained by the platform or vehicle on which they are developed. In this context, this paper makes a contribution to the state of the art with TE-NeXt. This network is a customised and efficient architecture for traversability estimation from sparse LiDAR point clouds based on a encoder-decoder topology that includes several modifications regarding: (i) the input features; (ii) the structure of encoder-decoder resolution levels; and (iii) the constitution of the 3D convolutional block. The modifications provide the robot with a deeper understanding of traversability in natural and unstructured environments. Thus, the designed architecture presents a trade-off solution, exhibiting superior performance in unstructured environments (the Rellis-3D dataset) while maintaining high reliability and robustness in urban environments (the SemanticKITTI dataset). The implementation of the presented approach is available at an open repository to ensure result reproducibility.
keywords:
Traversability Estimation , Artificial Intelligence , Computer Vision , Mobile Robotics, Autonomous Navigation
journal: Expert Systems with Applications
[…]
organization=University Institute for Engineering Research, Miguel Hernández University.,addressline=Avda. de la Universidad s/n, city=Elche (Alicante), postcode=03202, country=Spain
[…]
organization=Valencian Graduate School and Research Network of Artificial Intelligence (valgrAI),addressline=Camí de Vera s/n, city=Valencia, postcode=46022, country=Spain
[…]
organization=Univ. of Coimbra, ADAI, Department of Mechanical Engineering,addressline=Rua Luís Reis Santos, Pólo II, city=Coimbra, postcode=3030-289, country=Portugal
1 Introduction
@@
To enhance the efficacy of specialised techniques for off-road traversability estimation, this study proposes TE-NeXt, a LiDAR-based 3D sparse convolutional architecture capable of discerning traversable and non-traversable areas captured by a LiDAR sensor, enriching the perception of the environment with semantic information in addition to geometric information. In summary, the main contributions of this paper are:
  • 1.
@@
  • 3.
    The applicability of the proposed method in an autonomous navigation system in dynamic environments.
@@
The remaining sections of the paper are structured as follows: Section 2 outlines the prevailing methodologies employed in addressing the issue of traversability estimation. Section 3 presents the proposed architecture in detail. Following, Section 4 outlines the methodology employed to obtain the final architecture presented. Section 5 compares the proposed method with other state-of-the-art methods, demonstrating the advantages of the former. Finally, Section 6 presents the conclusions drawn from the study.
2 Related Work
@@
2.1 Vision-based traversability estimation
@@
2.2 Geometric and hybrid methods for traversability estimation
@@
Classification into road and off-road categories.
@@
Table 1: Overview of related works.
2.3 Self-supervised traversability estimation
@@
2.4 Current proposal
This paper presents TE-NeXt, an encoder-decoder architecture optimised for traversability estimation in unstructured environments. The approach presented here can be classified as a supervised method, in accordance with the paradigm described in Subsection 2.2, as it is a supervised paradigm method. However, it employs local point clouds, a distinctive feature that sets it apart from other methods, such as those described in [23, 24], which utilise point cloud maps and preprocess the data. Moreover, in contrast to the techniques outlined in Subsection 2.1, TE-NeXt employs an optical sensor that is invariant to changes in light conditions (LiDAR sensor). Furthermore, the proposed methodology is designed to overcome the limitations associated with self-supervised learning approaches. Unlike existing methods, it does not require a specific platform, making it a more general and secure solution for all types of vehicles. Table 1 outlines the distinctive features of the proposed approach in comparison to existing, state-of-the-art methodologies.
@@ 2.4 Current proposal
3 TE-NeXt: AN ARCHITECTURE FOR TRAVERSABILITY ESTIMATION
@@
3.1 Network Architecture Overview
@@
3.2 TE-NeXt residual block
@@ 4.1 Datasets
4.2 Training and evaluation process
Given the considerable discrepancy in the quantity of labelled data available from urban and natural environments, being the latter the primary focus of this research, the objective was to achieve a more balanced distribution of data across these domains. In particular, 10,700 clouds of unstructured environments have been selected, belonging to four of the five sequences in Rellis-3D. Additionally, the initial four SemanticKITTI sequences have been selected, thus ensuring 11,100 clouds of structured environments. Moreover, the validation and testing process has been conducted with previously unseen sequences from both datasets to evaluate the neural network’s generalisation capabilities.
@@ 4.2 Training and evaluation process
Table 2: Hyperparameters selected in training process.
Table 2 specifies the hyperparameters and settings used to train the model from randomly initialized weights. The optimizer used AdamW with an initial Learning Rate (LR) of 5e-3 along to Cosine Annealing scheduler causes the LR to follow a cosine function over the same training epoch. It performs periodic “warm restarts” where the learning rate is reset back to the initial value after a certain number of batches. This simulates the restarting of the learning process using the weights from the previous run as a good starting point. The periodic restarts help the model escape from local minima and saddle points. Since the model has been trained for a binary segmentation task, the F1 score is a suitable metric for imbalanced datasets and considers both: precision (the proportion of true positive instances among all positive predictions) and recall (the proportion of true positive predictions among all actual positive instances). This guarantees that the model is not optimising for one metric at the expense of the other. Moreover, should the F1 score on the validation set cease to improve or begin to decline, training can be terminated prematurely to avert overfitting.
4.3 Ablation study
@@
Figure 3: Design process of the proposed architecture, TE-NeXt, starting from a base MinkU-Net34C. All experiments are detailed in Table 3.
4.3.1 Input data
@@
  • Features: In contrast to 2D convolutions, where the input and output tensors have a fixed number of feature dimensions, typically called channels (e.g. RGB channels for images), 3D sparse convolutions allow sparse tensors to have an arbitrary number of feature dimensions. Each non-zero element in the sparse tensor can have a feature vector of arbitrary length. Therefore, a study of the most discriminating features for estimating traversability was carried out. Among the most representative are those shown in Figure 3 (experiments I.4, I.5, and I.6, referred to in Table 3). Nevertheless, the vector of 1’s (initialising the feature associated with each point to a value of 1) was selected as the primary candidate for future modifications, as it is equally discriminant and, at the time of inference, represents an improvement in efficiency.
4.3.2 Global Design
  • Cardinality: This term was introduced by ResNext [47], and refers to the number of parallel branches within each block. Related to this, a number of different configurations have been assessed. The initial configuration employed was the MinkU-Net34 network, comprising (2, 3, 4, 6, 2, 2, 2, 2, 2) branches per block. In order to reduce the number of model parameters, it was decided to test the MinkU-Net18 and MinkU-Net14, experiments G.1 and G.2. The smaller versions, in particular MinkU-Net14, exhibit a notable reduction in the number of branches per convolutional block in comparison to MinkU-Net34. This leads to a decrease in the number of filters and, consequently, in the number of trainable parameters in the model. In conclusion, through a meticulous examination of smaller architectural configurations, highly efficient MinkU-Net configurations have been identified that strike a balance between model size and performance. This illustrates that it is frequently feasible to significantly reduce the complexity of neural networks without compromising their capacity to learn and generalise. While this modification does not lead to an significant enhancement in performance, it does contribute to greater efficiency, as will be demonstrated in Table 6.
  • Skip connections: Unlike the U-Net architecture [37], Te-Next presents a variation in the connections that are made between the respective parts of the encoder and decoder. In addition to the four skip connections that were previously implemented in the experiments, two and three connections corresponding to experiments G.3 and G.4 have been evaluated. Notwithstanding the fact that both exhibited a reduction of 2% in terms of the F1 metric, the more robust performance options, experiment G.2 (later mentioned as the alternative experimental branch) and experiment G.4 (a configuration comprising just two connections), were subjected to further investigation in subsequent experiments. This has been influenced by concepts from related research [23, 24].
@@ 4.3.2 Global Design
Table 3: Experimental modifications carried out during the design of the proposed architecture, for which performance is presented in Figure 3.
4.3.3 Micro Design
@@
  • Kernel size: Conventional CNNs typically utilise relatively small kernels. In contrast, Transformer-based architectures and CNN designs inspired by the Transformer often employ significantly larger kernels, with sizes beginning at 7x7 and beyond. This allows them to capture longer-range dependencies and broader context, contributing to their strong performance on various vision tasks. To explore the solutions proposed in Transformer-based architectures, a study of the convolution size has been conducted. Experiments M.3, M.4 and M.5, depicted in Figure 3 and detailed in Table 3, demonstrate that a larger convolution window size leads to superior outcomes compared to the conventional convolution window sizes typically employed in convolutional networks (2, 3 or 5). Consequently, this variation results in an increase of 0.9% in the F1 metric with respect to the previous optimal result.
@@ 5 Results
This section offers a comprehensive performance evaluation of the proposed method, comparing it with established semantic segmentation and traversability estimation models from the literature. To ensure a fair comparison, we focus on accessible semantic segmentation models and methods originally designed for traversability estimation. The analysis covers a diverse range of approaches, including multi-layer perceptrons, traditional 2D convolutional networks, and cutting-edge sparse convolutions for efficient 3D data processing. As emphasised throughout this paper, TE-NeXt is an architecture optimised for unstructured environments. However, it can also be regarded as a trade-off solution in urban environments. In order to facilitate a comprehensive and structured comparison, this analysis will be divided into two distinct Subsections, each of which will focus on a specific type of environment.
@@
Table 4 presents the results of several semantic segmentation methods for the estimation of traversability in urban environments, using the validation sequence of the SemanticKITTI dataset.
5.1 Performance in structured environments
@@
In the context of traversability estimation, our proposed solution, while effective, is outperformed by methods like SqueezeSegv3 and SalsaNext when assessed in a 2D domain using image projections of 3D data (rows 3 and 5 of Table 4 respectively). However, this 2D evaluation process introduces significant challenges and limitations. The projection from three-dimensional (3D) to two-dimensional (2D) space inevitably leads to information loss due to occlusions and depth ambiguities, where multiple points may map to a single pixel. When projecting segmentation results back to 3D space, issues such as one-to-many mapping and interpolation errors at object boundaries emerge, which result in a noticeable decrease in performance metrics [50, 51].
@@
Thus, these methods have also been evaluated in a 3D context in order to compare their performance under the same conditions as the other methods (rows 4 and 6 of Table 4 respectively). A decrease in their performance is observed, which confirms the limitations of 2D to 3D projections. To illustrate, SqueezeSegv3 demonstrates a decrease in its F1 score from 96.5% in 2D to 95.8% in 3D, whereas SalsaNext exhibits a reduction from 95.9% to 94.9%. Consequently, while these methods may show superior results in 2D evaluations, they may not be optimal for critical traversability tasks in 3D real-world environments. In contrast, TE-NeXt demonstrates a solid balance between accuracy, precision, and reliability, achieving an F1 score of 96.0% in structured environments. Thus, making it more suitable and robust for semantic segmentation in challenging, 3D real-world scenarios.
@@
The comparison presented in Table 4 has been performed by training the semantic segmentation models from scratch, with the exception of Cyl3D, PVKD and P-SVM, which were adapted to the traversability problem in [25] using the same approach. Furthermore, it has included additional traversability estimation methods, such as ProtoNet [52], MPTI [53] and PBG [20]. These metrics have been collected from the manuscripts of the aforementioned approaches, which have been subjected to the same testing conditions and code is not accessible.
@@
Table 4: Traversability estimation performance of semantic segmentation methods in urban environments.
5.2 Performance in natural environments
Table 5 presents the results of various semantic segmentation methods for traversability estimation in natural environments using the Rellis-3D dataset. The comparison presented in Table 5 is of particular relevance, given that natural environments represent the primary focus of the research. Unlike in urban environments, it is remarkable the good performance of the most classic method, PointNet, outperforming more modern methods in most of the metrics extracted.